silvernight269:

x^2+x+68=37x-3x^2-13 =>0

silvernight269:

And no, I cannot/ not supposed to use the quadratic formula.

jhonyy9:

what you need to do with ?

silvernight269:

Well so far I got 4x^2-36x+81

silvernight269:

To equal 0

jhonyy9:

exactly very nice job

silvernight269:

After that... I’m stuck

jhonyy9:

what is the question ? you need to solve it getting roots of this quadratic or what ?

silvernight269:

What’s x to make it equal to 0

jhonyy9:

this mean that you need solve it for roots

silvernight269:

I guess yea

jhonyy9:

do you know the way ?

silvernight269:

It’s either 2x+? And 2x-? Or 4x+/-? And x+/- ?

jhonyy9:

4x^2 -36x +81 = 0 suppose rewrite it : in this above wrote way

jhonyy9:

4x*x -4*9x +9*9 = ?

silvernight269:

??

jhonyy9:

any idea ?

silvernight269:

No

jhonyy9:

4x^2 -36x +81 = 4x^2 -18x -18x +81 = or you the discriminant method - do you know it ?

silvernight269:

No

jhonyy9:

or use

jhonyy9:

D=b^2 -4ac

silvernight269:

Nope

jhonyy9:

really ?

silvernight269:

Yea

jhonyy9:

x_1,2 = (-b +/- sqrt(D) )/2a = may be now ?

silvernight269:

Nope

jhonyy9:

@ThisGirlPretty any help ?

silvernight269:

Sadly no

jhonyy9:

without discriminant method

silvernight269:

It’s ok

jhonyy9:

(2x-9)(2x-9) = ?

jhonyy9:

do you understand it now ?

jhonyy9:

4x^2 -36x +81 is a square perfect

jhonyy9:

(2x-9)^2

jhonyy9:

how you solve it now ?

jhonyy9:

write 2x-9=0

jhonyy9:

2x=9 and ?

jhonyy9:

divide both sides by 2 so what will get ?

jhonyy9:

x=9/2 so this the root of this quadratic

jhonyy9:

do you understand it now ?

jhonyy9:

ok was my pleasure bye bye